[Hallc_running] [Revised Logentry] Hall C RC Update: July 17, 2025

kinney at jlab.org kinney at jlab.org
Thu Jul 17 09:15:02 EDT 2025


Logentry Text:
--
During day shift (7/16), elastic scattering data from the LH2 target was collected at both high rate and low rate settings; this was in conjunction with the accelerator pushing the current limits up. Currents of 50 muA were achieved for short (several minute) periods; running at 45 muA was relatively stable. Collection of background from the dummy was not attempted in order to move to DIS in both spectrometers to allow setup of the ELREAL trigger. 

Dave and Bill set up the ELREAL thresholds during swing shift carried and then production DIS data was taken with the ELREAL levels set (both LH2 and Dummy). This was completed around 8pm at which time we moved to the x=0.25, Q2=3.3, z=0.5, theta_pq = 5.2 deg kinematic. Moving to the theta_pq=2.0 deg setting requires a crew to watch the spectrometer as it moves past 8.0 deg to 7.87 deg, hence this will take place Thursday during the maintenance day. 

We were unable to get the NMR to lock on the HMS dipole. Ellen tried to do this remotely, but was unable, hence the SIDIS runs have at present a poorly determined field. Before changing dipole current, Ellen will attempt to get the NMR to lock. In addition, we note that the coincidence peak is a bit to one side of the range, so Bill H will adjust to center it. Hence, we should probably repeat these runs later.

Beam returned around midnight, and we started collection of SIDIS data, starting off with a rates scan of all the targets, using 45 muA max. Before turn off at 7am, we were able to collect more LH2 data.

The plan for the day shift is to move the SHMS to 7.87 deg and determine the present field in the HMS, then move to the new field for z=0.36 running at theta_pq=2.0 deg. The hall is at present in controlled access, but there may be a need to go to restricted later. If anyone needs to do work in the hall, please contact me before making an access.

Update: A larger tour this afternoon will require going to restricted access.

There will be an RC meeting today at 2:30 pm in the meeting room on the 2nd floor of the Counting House. The usual Zoom information is below:

https://jlab-org.zoomgov.com/j/1618227115?pwd=8JkhFg6YNesVc3HNgWjwUSIKIu...
Meeting ID: 161 822 7115
Passcode: 253463

---

This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4420024
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20250717/70d3a3fb/attachment.htm>


More information about the Hallc_running mailing list